Find to three significant digits the charge and the mass of the following particles. Suggestion: Begin by looking up the mass of a neutral atom on the periodic table of elements in Appendix C of the textbook.
Question 8 1 pts a doubly ionized calcium atom, Cain kilograms] Question 9 1 pts the center of an ammonium molecule, modeled as an N3 ion [in coulombs Question 10 1 pts the center of an ammonium molecule, modeled as an N3- on in kilograms DQuestion 11 1 pts quadruply ionized nitrogen atoms, N4+, found in a hot star [in coulombs]
DQuestion 12 1 pts quadruply ionized nitrogen atoms N4+, found in a hot star lin kilograms] D Question 13 1 pts the nucleus of a nitrogen atom [in coulombs Question 14 1 pts the nucleus of a nitrogen atom [in kilograms] DQuestion 15 1 pts the molecular ion H2O [in coulombs
Question 16 1 pts the molecular ion H2O in kilograms]
Answer #1

8) for doubly ionized calcium atom, Ca++ = Ca2+ :

       mass = mass of calcium atom - (2)(mass of electron)

               = (40)(1.67*10-27 kg) - (2)(9.11*10-31 kg)

               = 6.68*10-26 kg

               = (6.68*10-26)(1/1.66*10-27) u

               = 40.2 u

9)

ammonia molecule (N3-) :

   charge = (3)(charge of electron)

              = (3)(-1.6*10-19 C)

              = -4.8*10-19 C

              = -480 zC

10) mass = mass of N atom + (3)(mass of electron)

             = (14)(1.67*10-27 kg) + (3)(9.11*10-31 kg)

            = 2.34*10-26 kg

            = 14 u

11)

for N4+ :

    charge = charge of four protons

               = (4)(1.6*10-19 C)

                =6.4*10-19 C

               = 640 zC

12) mass = mass of nitrogen atom - (4)(mass of electron)

            = (14)(1.67*10-27 kg) - (4)(9.11*10-31 kg)

            = 2.34*10-26 kg

            = 14 u

13)

for nucleus of a nitrogen atom (N77) :

it has seven protons and seven neutrons.

           charge = (7)(charge of proton)

                      = (7)(1.6*10-19 C)

                      = 11.2*10-19 C

                      = 1.12*10-18 C

                     = 1.12 aC

14) mass = mass of the nitrogen atom - (7)(mass of electron)

                    = (14)(1.67*10-27 kg) - (7)(9.11*10-31 kg)

                    = 2.34*10-26 kg

                    = 14 u

15)

for molecular ion H2O- :

         charge = charge of electron

                    = -1.6*10-19 C

                    = -160 zC

16) mass = mass of H2O molecule + (1)(mass of electron)

                  = (18)(1.67*10-27 kg) + (1)(9.11*10-31 kg)

                 = 3.01*10-26 kg

                 = 18.1 u